• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Aniol

    Log entry time 23:55:35 on August28,2009

    Entry number 286686

    keyword=shift summary

    16:00 no beam at start of shift, cryotarget was warming up at 1530 and is cooling down now
    18:00 target beginning to stabilize in pressure, temperature, heater power

    cryotarget operators are to read new instructions posted by poltarac
    by Dave Meekins

    19:00 asked for beam for Moller run, Sasha as notified

    21:00 0.5uA beam available now for Moller run

    23:55 Moller run completed